OpenStudy (anonymous):

f(x)= 3root 2x-1 find f(5)

jimthompson5910 (jim_thompson5910):

\[f(x)=3\sqrt{2x-1}\] \[f(5)=3\sqrt{2(5)-1}\] \[f(5)=3\sqrt{10-1}\] \[f(5)=3\sqrt{9}\] \[f(5)=3(3)\] \[f(5)=9\]

OpenStudy (anonymous):

Jim you are a genius.